OpenStudy (anonymous):

A sequence is defined by the recursive function f(n + 1) = f(n) – 2. If f(1) = 10, what is f(3)? a. 1 b. 6 c. 8 d. 30

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it would be B because your 1 is 10 and 3 would be 10-2-2 which is 6

OpenStudy (anonymous):

okay thank u @Musicdude

jimthompson5910 (jim_thompson5910):

Step 1) Replace every n with 1 f(n + 1) = f(n) – 2 f(1 + 1) = f(1) – 2 f(2) = 10 – 2 ... replaced f(1) with 10 f(2) = 8 Step 2) Repeat step 1 but instead of replacing n with 1, replace n with 2 Keep repeating these steps until you get to f(3)